OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7.Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7.

OpenStudy (anonymous):

@mathstudent55 @Compassionate

OpenStudy (compassionate):

Sorry, not good with upper maths!

OpenStudy (anonymous):

Thanks though! @Compassionate

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@YanaSidlinskiy @dan815 @just_one_last_goodbye @sleepyjess

OpenStudy (sleepyjess):

Sorry I learned this last year and do not remember anything about it.

OpenStudy (anonymous):

I have an example but I don't understand it either

OpenStudy (anonymous):

Example We can also write the equation of a parabola if we are given information about it. For example, let's write the equation of a parabola whose focus is at (-4, 0) and whose directrix is the line x = 8 We begin by sketching the given information and noting that our parabola has a horizontal axis and must open to the left. It therefore has the form (y - k )2 = 4p(x - h), p ≠ 0 We must find the values of h, k, and p. Since the distance from focus to directrix is 2p ( the distance from each to the vertex is p), 2p = 12 and |p| = 6. For a left-opening parabola, p = -6. This puts the vertex at the point (2, 0) with h = 2, and k = 0. The equation is therefore (y – 0)2 = -24 ( x – 2). This can be written as either y2 = –24 (x – 2) or x equals negative one divided by twenty four y squared plus two.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

I am so sorry, but I really don't get it either. \(\huge\ddot\frown\)

OpenStudy (yanasidlinskiy):

Sure, I can help! That's why we're here for!:) the focus is at (7, 0), the directrix in a parabola is equidistant from the vertex to the focus that is, the same distance from the vertex to the focus, is the distance from the vertex to the directrix so the directrix is at x= -7 notice the focus is to the right side of the origin and the directrix is to the left side that means the parabola is opening horizontally half-way between the directrix and the focus is, of course the vertex between x = 7 and x =-7, the half-way is the origin, (0, 0) so the center is at (0, 0) so the parabola looks like this: